A certain shade of light blue paint is made by mixing `1\frac{1}{2}` quarts of blue paint with 5 quarts of white paint. How much white paint would you need to mix with 4 quarts of blue paint?
1
Expert's answer
2020-11-09T19:50:48-0500

Let "x=" the amount of white paint mixing with 4 quarts of blue paint.

We can make a proportion


"\dfrac{1{1 \over 2}}{5}=\dfrac{4}{x}"

Solve for "x"


"\dfrac{3}{2}x=4(5)"

"x=20(\dfrac{2}{3})"

"x=\dfrac{40}{3}"

"x=13{1 \over 3}"

We need to mix "13{\dfrac{1}{3}}" quarts of white paint with 4 quarts of blue paint.